OpenStudy (anonymous):

A ladder 20 feet long leans against a wall and forms an angle of 30° with the ground. How far from the wall is the ladder?

OpenStudy (anonymous):

\[10\sqrt{3}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

we know that \[\cos = \frac{ adjacent }{ hypothenuse }\] here, \[\cos(30)=\frac{ ? }{ 20 }\]